OpenStudy (anonymous):

Which of the following is a solution to the system y=-3x+5 and y=2x+5? A.(1,7) B.(5,0) C.(2,-1) D.(0,5)

OpenStudy (anonymous):

\[y=-3x+5\] \[y = 2x+5\] so \[-3x+5 = 2x+5\] -3x = 2x so x must be 0 now y = -3(0) + 5 = 5

OpenStudy (mathmagician):

subtract one equation from another, and you get (y-y)=-3x+5-2x-5, so x is equal to 0. if you use value x=0 in first equation, you get y=-3*0+5, so answer is (0;5)